# SVGFusion: Scalable Text-to-SVG Generation via Vector Space Diffusion

### v1

@@ Abstract
The generation of Scalable Vector Graphics (SVG) assets from textual data remains a significant challenge, largely due to the scarcity of high-quality vector datasets and the limitations in scalable vector representations required for modeling intricate graphic distributions. This work introduces SVGFusion, a Text-to-SVG model capable of scaling to real-world SVG data without reliance on a text-based discrete language model or prolonged SDS optimization. The essence of SVGFusion is to learn a continuous latent space for vector graphics with a popular Text-to-Image framework. Specifically, SVGFusion consists of two modules: a Vector-Pixel Fusion Variational Autoencoder (VP-VAE) and a Vector Space Diffusion Transformer (VS-DiT). VP-VAE takes both the SVGs and corresponding rasterizations as inputs and learns a continuous latent space, whereas VS-DiT learns to generate a latent code within this space based on the text prompt. Based on VP-VAE, a novel rendering sequence modeling strategy is proposed to enable the latent space to embed the knowledge of construction logics in SVGs. This empowers the model to achieve human-like design capabilities in vector graphics, while systematically preventing occlusion in complex graphic compositions. Moreover, our SVGFusion’s ability can be continuously improved by leveraging the scalability of the VS-DiT by adding more VS-DiT blocks. A large-scale SVG dataset is collected to evaluate the effectiveness of our proposed method. Extensive experimentation has confirmed the superiority of our SVGFusion over existing SVG generation methods, achieving enhanced quality and generalizability, thereby establishing a novel framework for SVG content creation. Code, model, and data will be released at: https://ximinng.github.io/SVGFusionProject/
Figure 1: Example SVGs generated by our SVGFusion. Noteworthy characteristics of the SVGs generated by our new method include: (a) primitive ordering aligned with human design principles, (b) a clear and systematic layering structure of vector primitives, and (c) high editability.
11footnotetext: Corresponding author.
@@ 1 Introduction
Scalable Vector Graphics (SVGs) have become essential in modern digital design, utilizing mathematical definitions to represent shapes such as paths, curves, and polygons instead of relying on pixel-based raster graphics. This vector-based approach enables SVGs to scale across resolutions without any loss of detail, offering precise quality control. Additionally, SVGs are efficient in terms of storage and transmission due to their compact structure, and they provide high levels of editability, allowing designers to make fine adjustments to graphic elements. As a result, SVGs play an essential role in applications like web design and user interfaces, including icons, logos, and emojis.
@@
The task of SVG generation has attracted increasing attention in recent years. Existing methods can be categorized into two types: optimization-based and language model-based. Optimization-based methods [8, 52, 56, 34, 46, 18, 63, 64] generally initialize an SVG in vector space and then rasterize it into pixel space via differentiable rendering [19] for loss computation. Gradients are back-propagated to the vector space to update the parameters of the SVG. Although optimization-based methods can produce good results by leveraging priors from large pre-trained models like CLIP [34] or Stable Diffusion [38], they are very time-consuming. Moreover, limited by the use of differentiable rasterizers [19], these methods [18, 63, 64, 65] only consider a few types of ‘differentiable’ commands in SVGs, e.g., Cubic Bézier curves. Furthermore, the results generated by these methods are difficult to edit as the primitives often intertwine or occlude each other.
@@
Alternatively, language model-based methods [11, 3, 58, 59, 61, 50] treat SVG generation as a sequential task and employ a language model, like RNNs or Transformers, to predict in an auto-regressive manner. However, these methods can become inefficient when the command sequence of an SVG is too long. Additionally, it is challenging for these models to learn the complex syntax and numerous primitive types without extensive data. Consequently, these methods often simplify the task by limiting SVG the representation to one primitive type, <path>, and three path commands: Move to(M), Line to(L), and Cubic Bézier(C). This simplification mainly restricts their output to generating black-and-white icons and single-letter shapes, thus limiting their applicability in real-world SVG applications. Moreover, existing methods generally overlook the inherent creation logic of an SVG, leading to unreasonable constructions.
In this work, we propose SVGFusion, a novel model designed for text-based SVG generation. This model can output SVGs that are decomposable, editable, and reflect human-like design logic, as demonstrated in Fig. 1. Inspired by the recent success of Latent Diffusion Models (LDMs) [38] in the Text-to-Image (T2I) task, our SVGFusion adopts an LDM-like framework wherein a continuous vector space is learned to generate SVGs. Specifically, SVGFusion comprises two core modules: a Vector-Pixel Fusion Variational Autoencoder (VP-VAE) and a Vector Space Diffusion Transformer (VS-DiT). The VP-VAE is tasked with learning a latent space for vector sketches, whereas the VS-DiT is responsible for generating a latent code within this space, conditioned on the input text prompt. This architecture enables our model to capitalize on the successes observed in the T2I tasks. The VP-VAE employs a transformer-based architecture, which enhances SVGFusion’s capability to effectively model the latent space and supports a broader range of primitives. Additionally, SVGFusion can continuously enhance its interpretive and generative capacities by leveraging the scalability of the VS-DiT, which can be expanded by adding more VS-DiT blocks.
More importantly, SVGFusion is capable of emulating human design principles in SVGs, ensuring more reasonable SVG construction and improved editability. To achieve this, we introduce a rendering sequence modeling strategy, where a batch learning unit is structured according to the SVG rendering sequence to simulate human design processes (as shown in Fig. 7). Furthermore, we incorporate visual data—that is, the rasterizations of the corresponding SVGs—and feed them along with the SVGs into our VP-VAE to learn a latent space that embeds human design principles. To simulate real-world SVG synthesis and evaluate our proposed method, we collected approximately 240,000 high-quality, human-designed SVGs, including color emojis and icons. We also established an automated processing pipeline for lossless data refinement. Extensive experiments demonstrate that SVGFusion not only enhances the visual quality of SVGs but also exhibits strong generalizability.
The contribution of this work is three-fold:
@@
  • A Novel, Efficient, and Scalable SVG Generation Model: We introduce SVGFusion, which adopts a popular Text-to-Image framework for Text-to-SVG generation. This model is not only efficient and scalable but also supports a broader range of SVG primitives.
@@
  • Specific Module Design and Training Strategy to Learn Creation Logic: To capture the creation logic of man-made SVGs and ensure reasonable SVG construction, we propose a novel Vector-Pixel Fusion Variational Autoencoder (VP-VAE) that processes both SVGs and their corresponding rasterizations. Additionally, a novel rendering sequence modeling strategy is implemented to prepare the training data. To the best of our knowledge, this is the first Text-to-SVG model that explicitly considers the creation logic of SVGs.
@@
  • A Large-Scale, High-Quality SVG Dataset and Comprehensive Evaluation: We have curated a dataset of approximately 240k high-quality, human-designed SVGs and conducted extensive experiments to validate SVGFusion’s effectiveness in SVG generation, thereby setting a new benchmark for this task.
@@ 2.1 Vector Graphics Generation
Scalable Vector Graphics (SVGs) are widely used in design due to their advantages, such as geometric manipulability, resolution independence, and compact file structure. One approach to SVG generation involves training neural networks to output predefined SVG commands and attributes [11, 21, 3, 36, 61, 50]. These networks typically use architectures like RNNs [11, 36], VAEs [21], and Transformers [3, 61, 50]. However, the lack of large-scale vector datasets limits their ability to generalize and create intricate graphics.
Currently, the vector graphics field lacks datasets comparable in scale to ImageNet [6], with most datasets focused on narrow areas such as monochromatic vector icons [5], emojis [9], and fonts [58]. An alternative to directly training an SVG generation network is optimizing to match a target image during the evaluation phase.
Li et al.[19] introduced a differentiable rasterizer to bridge vector graphics and raster images. Recent work has used differentiable rasterizers to overcome dataset limitations[52, 24, 46, 63, 16, 65]. This method optimizes SVG parameters based on a pretrained vision-language model. Advances in models like CLIP [34] have enabled successful sketch generation methods, such as CLIPDraw [8] and CLIPasso [56], while DreamFusion [33] has shown diffusion models’ superior generation abilities. VectorFusion [18], DiffSketcher [63], and SVGDreamer [64] combine differentiable rasterizers with text-to-image diffusion models to generate vector graphics, achieving promising results in iconography and sketching. However, these methods still suffer from limited editability and graphical quality. Recent studies [51, 65] have combined optimization-based approaches with neural networks to improve vector representations by incorporating geometric constraints. We propose SVGFusion, an innovative model for scalable SVG generation that operates in a continuous vector space, overcoming the limitations of discrete language models and intensive optimization techniques to generate high-quality SVGs at scale.
@@ 2.2 Diffusion Model
Figure 2: An overview of our SVGFusion pipeline. (a) Our pipeline begins with the neural representation of SVGs, where XML-defined SVG tensors are transformed into a learnable matrix to derive an SVG embedding (Sec.3.1). (b) We propose the Vector-Pixel Fusion Variational Autoencoder (VP-VAE, Sec.3.2) within a transformer-based architecture to encode vector embeddings alongside pixel-level features into a latent vector space. The resulting vectors are subsequently decoded using a transformer decoder, which parallels the encoder, to reconstruct vector graphics. (c) The Vector Space Diffusion Transformer (VS-DiT, Sec.3.3) is then trained within the latent space constructed by the VP-VAE. Textual features extracted from the text prompt using the CLIP[34] models are incorporated into each VS-DiT block. The generative capability of SVGFusion can be continuously enhanced by stacking additional VS-DiT blocks.
@@ 3 Methods
General-purpose SVGs, often designed manually, are composed of numerous vector primitives, which creates substantial challenges for language models when modeling sequences directly in the SVG representation space due to increased computational demands. Drawing inspiration from recent advancements in T2I methods [38, 31] developed within latent spaces, this strategy facilitates the construction of longer sequences and the application of more intricate models.
As shown in Fig. 2, our SVGFusion pipeline proceeds through three main stages. (a) The pipeline begins by encoding SVGs into a neural representation. Here, XML-defined SVGs are converted into a learnable matrix to yield an SVG embedding. To introduce human design priors, we propose a rendering sequence modeling approach and construct pairs of SVG codes and their corresponding renderings. (b) We introduce the Vector-Pixel Fusion Variational Autoencoder (VP-VAE), a Transformer-based architecture that encodes both vector embeddings and pixel-level features into a latent vector space. The encoded vectors are subsequently passed through a Transformer decoder, which parallels the encoder structure, to reconstruct vector graphics. (c) Finally, the Vector Space Diffusion Transformer (VS-DiT) is trained within the latent space established by the VP-VAE. Textual information is extracted from text prompts using the CLIP text encoder [34] and is incorporated into each VS-DiT block. Operating within this latent space, SVGFusion allows for flexible model scalability by enabling additional VS-DiT blocks to be stacked, thereby refining the generative performance.
Let an SVG file be defined as a collection of $N$ vector primitives: SVG $\mathcal{G}=\{\mathrm{E}_{1},\mathrm{E}_{2},\dots,\mathrm{E}_{n}\}=\{\mathrm{E}_{i}\}_{i=1}^{N}$, where each $\mathrm{E}_{i}$ represents an element-level primitive, such as <path>, <rect>, <circle>, etc. $\mathrm{E}_{i}=\{\mathrm{S}_{i,j},\mathrm{F}_{i,j},\alpha_{i,j}\}_{j=1}^{M_{i}}$, where $\mathrm{S}_{i,j}$ is the $j$-th command-level primitive in the $i$-th element, $\mathrm{F}_{i}\in\{r,g,b\}$ is the $i$-th element color property and $\alpha_{i}\in\{0,1\}$ is the visibility of the $i$-th element. $M_{i}$ indicates the total number of commands in $\mathrm{E}_{i}$. The element <path> may consist of several commands, while other elements consist of only one command.
@@ 3.1 SVG Representation
Vector graphics consist of machine instructions composed of a series of XML elements (e.g. <path> or <rect>) and attributes (e.g. d, rx or ry), which we tensor quantize into a learnable matrix. Inspired by previous works [3, 58], we define an SVG in matrix form so that it can be easily represented by tensor quantization. While previous works [3, 58, 61] simplify the SVG modeling task by limiting the command set, thereby constraining the generalizability of SVG representations, our work broadens the SVG element and command sets to accommodate a more extensive range of vector forms. This enhancement significantly improves the model’s ability to learn from real-world data.
SVG Embedding. In [3, 61, 50, 58], only a single SVG element, <path>, is employed to approximate all other SVG elements, which inherently leads to information loss and reduces fidelity to human-created SVGs. To address this limitation, we extend the SVG element set beyond <path> to encompass multiple distinct elements (see Supplementary Sec. C), thereby enhancing the model’s adaptability and precision in representing complex SVG designs. We first transform the SVG code into an SVG matrix by converting each primitive to an individual vector with rules: $\vec{E}_{i}=\textstyle\sum_{j=1}^{M_{i}}\mathrm{S}_{i,j}=(\rho,\tau,\mu_{0},\nu_{0},\mu_{1},\nu_{1},\mu_{2},\nu_{2},\mu_{3},\nu_{3})_{j}^{i}$, where $\rho$ and $\tau$ represent the element and command types, respectively. If $\rho$ is the <path> element, then $(\mu_{0},\nu_{0})$ and $(\mu_{3},\nu_{3})$ represent the beginning and end points, while $(\mu_{1},\nu_{1})$ and $(\mu_{2},\nu_{2})$ are control points. Then, to establish interconnections among the adjacent commands, we set the end point of command $(\mu_{3},\nu_{3})$ equal to the beginning point $(\mu_{0},\nu_{0})$ of the subsequent command in each individual path. If $\rho$ is <rect>, then $(\mu_{i},\nu_{j})_{i=0,j=0}^{3,3}$ correspond to the width, height, x, y, rx, and ry, respectively. We decompose each SVG primitive into basic commands, combining them into a structured matrix form. This results in a structured SVG matrix $\mathcal{G}\in\mathbb{R}^{(\sum_{i}^{N}M_{i})\times 10}$, where $\sum_{i}^{N}M_{i}$ indicates individual basic commands.
Rendering Sequence Modeling. Our goal is to develop an SVG synthesis method that aligns with human design logic and enables the model to learn SVG creation through imitative learning, by closely simulating the sequence in which humans draw vector graphic primitives. Specifically, we guide the model to construct SVGs incrementally, following the drawing order of primitives in real data, allowing it to observe progressive changes in the SVG structure across batch dimensions. In addition, we inject visual priors into the process. We use pre-trained DINOv2 [30] $f_{\mathrm{dino}}(\cdot)$ to extract visual features from each SVG rendering and overlay them to form visual prior variations. Such design enables the model to learn the SVG drawing process through both visual and geometric changes. To enhance the model’s comprehension of sequential information, we incorporate rotary position embeddings (RoPE) [48], similar to those used in advanced large language models [28, 53], to capture dependencies across different sequence positions.

### v2

@@ Abstract
In this work, we introduce SVGFusion, a Text-to-SVG model capable of scaling to real-world SVG data without relying on text-based discrete language models or prolonged Score Distillation Sampling (SDS) optimization. The core idea of SVGFusion is to utilize a popular Text-to-Image framework to learn a continuous latent space for vector graphics. Specifically, SVGFusion comprises two key modules: a Vector-Pixel Fusion Variational Autoencoder (VP-VAE) and a Vector Space Diffusion Transformer (VS-DiT). The VP-VAE processes both SVG codes and their corresponding rasterizations to learn a continuous latent space, while the VS-DiT generates latent codes within this space based on the input text prompt. Building on the VP-VAE, we propose a novel rendering sequence modeling strategy which enables the learned latent space to capture the inherent creation logic of SVGs. This allows the model to generate SVGs with higher visual quality and more logical construction, while systematically avoiding occlusion in complex graphic compositions. Additionally, the scalability of SVGFusion can be continuously enhanced by adding more VS-DiT blocks. To effectively train and evaluate SVGFusion, we construct SVGX-Dataset, a large-scale, high-quality SVG dataset that addresses the scarcity of high-quality vector data. Extensive experiments demonstrate the superiority of SVGFusion over existing SVG generation methods, establishing a new framework for SVG content creation.
Figure 1: Example SVGs generated by our SVGFusion. Our proposed method, SVGFusion can generate SVGs with (a) reasonable construction, (b) a clear and systematic layering structure, and (c) highly editability.
@@ 1 Introduction
Scalable Vector Graphics (SVGs) are indispensable in modern digital design due to their ability to represent shapes—such as paths, curves, and polygons—using mathematical definitions instead of pixel-based raster graphics. This vector-based approach enables SVGs to scale seamlessly across different resolutions without any loss of detail, ensuring precise quality control. Additionally, SVGs are highly efficient in terms of storage and transmission due to their compact structure. They also offer exceptional editability, allowing designers to make fine adjustments to graphic elements. Consequently, SVGs play a critical role in applications such as web design, user interfaces, and the creation of icons, logos, and emojis.
@@
The task of SVG generation has garnered increasing attention in recent years. Existing methods can be broadly categorized into two types: optimization-based and language-model-based. Optimization-based methods, such as those described in [9, 52, 56, 35, 46, 20, 65, 68], typically initialize an SVG in vector space and then rasterize it into pixel space using differentiable rasterizers [21] for loss computation. Gradients are back-propagated to the vector space to update the SVG parameters. While these methods can produce high-quality results by leveraging priors from large pre-trained models such as CLIP [35] or Stable Diffusion [38], they are computationally expensive and often support only a limited set of differentiable SVG commands, such as Cubic Bézier curves. Moreover, the generated results are often challenging to edit, as the primitives tend to intertwine or occlude one another.
@@
On the other hand, language-model-based methods [13, 4, 58, 59, 63, 49, 66] treat SVG generation as a sequential task, using language models like RNNs or Transformers to predict commands in an auto-regressive manner. However, these methods struggle with efficiency when the sequence of an SVG becomes excessively long and face challenges in learning the complex syntax and diverse primitive types of SVGs without extensive training data. Consequently, these methods often restrict their output primarily to black-and-white icons and single-letter shapes, significantly limiting their applicability in real-world SVG applications. Moreover, the sequential nature of these models can lead to error accumulation and limited global coherence in the generated outputs.
In this work, we propose SVGFusion, a novel model designed for text-based SVG generation. SVGFusion is capable of producing high-quality SVGs that are both decomposable and editable, as illustrated in Fig. 1. Inspired by the recent success of Latent Diffusion Models (LDMs) [38] in the Text-to-Image (T2I) task, SVGFusion adopts an LDM-like framework to learn a continuous latent space for SVG generation. It comprises two main components: 1) Vector-Pixel Fusion Variational Autoencoder (VP-VAE): A transformer-based VAE that learns a unified latent space for vector graphics by combining both SVG code and its rasterized image counterpart. 2) Vector Space Diffusion Transformer (VS-DiT): A scalable diffusion backbone operating in the learned latent space, conditioned on text prompts to generate highly-editable SVGs.
Our model introduces three key designs. (1) Vector-Pixel Fusion Encoding. To learn a more representative latent space, features from both SVG code and its rasterization are used, leading to SVGs with higher visual quality. During the learning phase for vector graphics, the VP-VAE encoder processes both the SVG code and its rasterization as inputs, while the decoder reconstructs the SVG code. Before being fed into the encoder, the SVG code is converted into a matrix representation with geometric attributes, colors, and opacity. Additionally, a pre-trained DINOv2 [31] is employed to extract visual features from the SVG rendering. (2) Rendering Sequence Modeling. Existing methods often struggle to generate editable SVGs because they lack an understanding of how SVG codes correspond to their visual representations. To address this challenge, we introduce a rendering sequence modeling strategy that presents the model with an incremental accumulation of SVG codes and their corresponding renderings. This strategy enables the model to learn the design logic of human creators, leading to SVGs with more logical constructions and enhanced visual quality, as shown in Fig. 1(a). (3) Incorporation of DiT Blocks. Motivated by the success of DiT [32] in Text-to-Image (T2I) tasks and their scalability, we incorporate VS-DiT blocks and perform the diffusion process in the latent space specifically learned for SVGs. This design allows SVGFusion’s generative capabilities to be continuously enhanced by scaling the VS-DiT through the addition of more blocks.
To train and evaluate our model, we curated a dataset of approximately 240,000 high-quality, human-designed SVGs from different sources. Additionally, we designed an automated data pre-processing pipeline to trim the SVGs losslessly. The contributions of this work are threefold:
@@
  • A Novel and Scalable SVG Generation Model: We introduce SVGFusion, which adapts a popular Text-to-Image framework for Text-to-SVG generation. This model is scalable and capable of producing high-quality SVGs that are highly editable.
@@
  • Specific Module Design and Training Strategy for SVGs: We propose a Vector-Pixel Fusion Variational Autoencoder (VP-VAE) that learns the latent space for SVGs from both SVG codes and their corresponding rasterizations. A novel rendering sequence modeling strategy is also designed to enable the model to understand the creation logic of SVGs, thereby enhancing the visual quality and editability of the generated outputs.
@@
  • A Large-Scale, High-Quality SVG Dataset and Comprehensive Evaluation: We construct SVGX-Dataset, a collection of approximately 240,000 high-quality, human-designed SVGs, and conduct extensive experiments to validate SVGFusion’s effectiveness in SVG generation, establishing a new benchmark for this task.
@@ 2.1 Vector Graphics Generation
Scalable Vector Graphics (SVGs) are widely used in design due to their key advantages, including geometric manipulability, resolution independence, and compact file sizes. One approach to SVG generation involves training neural networks to output predefined SVG commands and attributes [13, 23, 4, 37, 63, 49]. These networks typically use architectures like RNNs [13, 37], VAEs [23], and Transformers [4, 63, 49]. However, their ability to generate intricate and diverse vector graphics is constrained by the lack of large-scale datasets for training.
Unlike raster image generation, which benefits from large-scale datasets like ImageNet [7], the vector graphics domain has limited publicly available datasets, most of which focus on narrow domains such as monochrome icons [6], emojis [10], and fonts [58]. As an alternative to directly training an SVG generation network, optimization-based approaches have been explored, where vector parameters are iteratively refined to match a target image during inference.
Li et al. [21] introduced DiffVG, a differentiable rasterizer that bridges vector and raster domains, enabling gradient-based SVG optimization. Subsequent works [52, 26, 46, 65, 18, 69, 67] leveraged differentiable rasterizers with vision-language models, such as CLIP [35], to enhance text-guided sketch generation [9, 56]. Meanwhile, diffusion models like DreamFusion [34] demonstrated superior generative capabilities. Recent methods, including VectorFusion [20], DiffSketcher [65], and SVGDreamer [68], integrate differentiable rasterization with diffusion models, achieving promising results in sketching and iconography. However, they still suffer from limited editability, redundant geometry, and suboptimal graphical quality. To address these issues, hybrid approaches [51, 69] incorporate geometric constraints to refine SVG paths, but remain restricted to SDS-optimized structures and path-based primitives. More recently, VecFusion [50] introduced an image-conditioned diffusion model for vector font synthesis, though its scope is limited to fonts. In this work, we present SVGFusion, a novel model for scalable SVG generation in a continuous vector space. By moving beyond discrete language models and intensive optimization, SVGFusion enables high-quality, diverse, and editable SVG generation while overcoming key limitations of prior methods.
@@ 2.2 Diffusion Model
Figure 2: An Overview of SVGFusion. (a) The pipeline begins with the representation of SVGs, where XML-defined SVG code is converted into an SVG embedding (Sec. 3.1). (b) We first train a Vector-Pixel Fusion Variational Autoencoder (VP-VAE, Sec. 3.2) with a transformer-based architecture to learn a continuous latent space for SVGs by incorporating features from both SVG codes and their rendered images. (c) The Vector Space Diffusion Transformer (VS-DiT, Sec. 3.3) is then trained within the learned latent space to generate new latent codes conditioned on input text descriptions.
@@ 3 Methods
Our task is to generate editable SVGs from input text prompts. As illustrated in Fig. 2, our model first trains a Vector-Pixel Fusion Variational Autoencoder (VP-VAE) to learn a latent space $\mathcal{Z}$ for SVGs. Next, a Vector Space Diffusion Transformer (VS-DiT) is trained within this latent space to generate new latent codes conditioned on text prompts. Once trained, given an input text and a randomly sampled latent code, our model produces an SVG that is semantically aligned with the text. In the following sections, we first describe the process of converting SVG code into SVG embeddings (Sec.3.1), followed by discussions on VP-VAE (Sec.3.2) and VS-DiT (Sec. 3.3).
@@ 3.1 SVG Representation
Vector graphics consist of machine instructions composed of a series of XML elements (e.g. <path> or <rect>), commands (e.g. M, C in <path> element), and attributes (e.g. d, r or fill). Inspired by prior works [4, 58], we transform these instructions into a structured, rule-based matrix representation. We formally define an SVG code as a collection of $n$ vector primitives: SVG $\mathcal{G}=\{\mathrm{E}_{1},\mathrm{E}_{2},\dots,\mathrm{E}_{n}\}=\{\mathrm{E}_{i}\}_{i=1}^{n}$, where each $\mathrm{E}_{i}$ corresponds to an SVG element such as <path>, <rect>, <circle>, etc. Each element is defined as: $\mathrm{E}_{i}=\{\mathrm{S}_{i,j},\mathrm{F}_{i,j},\alpha_{i,j}\}_{j=1}^{M_{i}}$, where $\mathrm{S}_{i,j}$ is the $j$-th command in the $i$-th element, $\mathrm{F}_{i}\in\{r,g,b\}$ and $\alpha_{i}\in[0,1]$ correspond to the color property and the visibility of the $i$-th element, respectively. $M_{i}$ indicates the total number of commands in $\mathrm{E}_{i}$. Thus, the total number of primitive commands in an SVG is $N=\sum_{i=1}^{n}M_{i}$. In our work, we set $N=1024$, determining the maximum number of commands in an SVG representation. Notably, the <path> element consists of multiple commands, while other elements typically contain only a single command.
Figure 3: Illustration of the SVG embedding process. SVG code is initially converted into a matrix representation that includes geometric attributes, colors, and opacity. This matrix is subsequently mapped into a tensor via SVG embeddings.
Figure 3 illustrates the process of converting an SVG code into an SVG embedding. We begin by transforming each primitive into a vector representation, which we then organize into an SVG matrix. Specifically, each command $\mathrm{S}_{i,j}$ is represented as $\mathrm{S}_{i,j}=(\rho,\tau,\mu_{0},\nu_{0},\mu_{1},\nu_{1},\mu_{2},\nu_{2},\mu_{3},\nu_{3})_{j}^{i}$, where $\rho$ and $\tau$ represent the type index of an element and command, respectively. Taking the element <rect> as an example, its element index $\rho$ is 6, and its command index $\tau$ is 0 (as it only has no <path> command), its $(\mu_{i},\nu_{j})_{i=0,j=0}^{3,3}$ correspond to x, y, rx, ry, width and height. Further details are provided in Table S1 in Supplementary. After converting into a matrix, we apply an embedding layer to transform $\rho$ and $\tau$ from discrete indices into continuous representations while normalizing coordinates and colors in the matrix. Finally, we obtain the SVG embedding by adding the positional embedding.
Compared to previous approaches [4, 58, 63], our method supports a broader range of SVG primitives, including more elements (e.g., <circle>, <rect>) and commands (e.g. Q, A for <path>). A full list of primitives supported by our model is shown in Table S1 of Supplementary. This enhancement significantly improves the model’s ability to learn from real-world data, making the generated SVGs more structured and editable.